One number is 25% of another. The larger number is 12 more than the smaller. What are the nunmbers?
Answer by Maths68(1474)

Let
Smaller number = x
Larger number = y


One number is 25% of another.
x=25%y
x=25/100(y)
x=1/4 (y)
x=y/4 ............(1)
The larger number is 12 more than the smaller
y=x+12
Subtitue the value of x in above equation
y=y/4+12
y=(y+48)/4
4y=y+48
4y-y=48
3y=48
y=48/3
y=16
Subtitue the value of y in (1)
x=16/4
x=4
numbers are 4 and 16